# Pretraining with random noise for uncertainty calibration

### v1

@@ Acknowledgements
J.C. conceived the project. J.C., and S.P. designed the model. J.C. performed the simulations. J.C. and S.P. analyzed the data. J.C. and S.P. wrote the manuscript and designed the figures.
@@
Supplementary Figures
Supplementary Fig. 1: Random noise for pretraining and data for subsequent training
(a) Example of random noise (32×32×3) used in the pretraining phase of the network. This noise serves as an input during the initial random noise pretraining stage. (b) Statistics of the random noise and its corresponding random labels. The random noise is sampled from a Gaussian distribution with zero mean and a standard deviation of one. The random labels are sampled from a uniform distribution. In each iteration, random noise and labels are resampled and are not paired. (c) CIFAR-10 50, a benchmark natural image dataset used in the subsequent data training phase. CIFAR-10 consists of ten object and animal categories: Airplane, Automobile, Bird, Cat, Deer, Dog, Frog, Horse, Ship, and Truck.
Supplementary Fig. 2: Learning curve in random noise pretraining
(a) The pretraining process, where neural networks are trained on random noise before encountering real data. (b-c) Learning curves during random noise pretraining: (b) Training loss (c) Training accuracy. Note that the accuracy remains at chance level, as there is no explicit correlation between the input (random noise) and the output labels.
Supplementary Fig. 3: Learning curve in subsequent data training stage
(a) After pretraining with random noise, the network is subsequently trained with real data. (b-e) Learning curve during random noise pretraining. (b) Training loss. (c) Training accuracy. (d) Test loss. (e) Test accuracy. In each learning curve, the orange line represents the network trained solely with data (without random noise pretraining), while the blue line represents the network pretrained with random noise.
Supplementary Fig. 4: Uncertainty calibration via random noise pretraining with biologically plausible learning
(a) Comparison of backpropagation 67 and feedback alignment 68. Backpropagation is the standard training method for artificial neural networks but is not considered biologically plausible due to the weight transport problem 69, 70, which requires a symmetric backward pathway for transmitting error signals. Feedback alignment, on the other hand, is a biologically plausible alternative that does not rely on weight transport, instead using a random matrix $\mathbf{B}$ to replace the backward synaptic pathway. (b) Test loss during random noise pretraining and subsequent data training. Blue indicates networks pretrained with random noise, and orange indicates networks trained solely with data (without random noise pretraining). (c) Test accuracy during subsequent data training. (d) Histogram of predictive confidence, with vertical lines indicating the averaged accuracy and confidence. (e) The difference between the averaged confidence and accuracy of predictions (w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$; w/o vs. zero, Wilcoxon signed-rank test, NS, $P<10^{-3}$; w/ vs. zero, Wilcoxon signed-rank test, NS, $P=0.752$). (f) Reliability diagram showing expected sample accuracy as a function of binned confidence. (g) Expected calibration error (w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$).
@@
Supplementary Fig. 5: Pre-calibration of uncertainty in toy-model network
@@
Supplementary Fig. 6: Pre-calibration of uncertainty in high-dimensional network

### v2

@@ Acknowledgements
Conception and design: J.C. Development of the methodology: J.C. and S.P. Funding acquisition: S.P. Collection of data: J.C. Formal analyses: J.C. and S.P. Writing the original draft: J.C. and S.P. Writing, reviewing and/or revising the manuscript: J.C. and S.P. Study supervision: S.P.
@@
Supplementary Materials
@@
Supplementary Fig. 1: Measuring confidence calibration of a model network through a reliability diagram.
(a) Sample images were presented to a trained neural network, and its predictions were recorded ($n_{\text{trial}}=10000$). Specifically, confidence and predicted answers were measured in each trial. (b) Confidence histogram. Each trial is binned according to its confidence level. (c) In each batch of predictions binned by confidence, the correct ratios were measured. The diagram shows two example batches with prediction trials at different confidence levels. In an ideally calibrated model, the correct ratios in each batch should match the confidence levels. (d) Reliability diagram. This diagram shows the model’s accuracy as a function of confidence levels for each batch. In an ideally calibrated model (gray bars), accuracy should align with confidence levels across all predictions, represented by a diagonal line.
Supplementary Fig. 2: Learning curve in random noise pretraining and subsequent data training stage.
(a) The pretraining process, where neural networks are trained on random noise before encountering real data. (b) After pretraining with random noise, the network is subsequently trained with real data. (c-d) Loss curves during random noise pretraining: (c) Training loss. (d) Test loss. (e-f) Loss curve during subsequent data training: (e) Training loss. (f) Test loss. (g-h) Accuracy curves during random noise pretraining: (g) Training accuracy. (h) Test accuracy. Note that the accuracy remains at chance level, as there is no explicit correlation between the input (random noise) and the output labels. (i-j) Accuracy curve during subsequent data training: (i) Training accuracy. (j) Test accuracy. In each learning curve, the orange line represents the network trained solely with data (without random noise pretraining), while the blue line represents the network pretrained with random noise.
Supplementary Fig. 3: Uncertainty calibration via random noise pretraining with biologically plausible learning.
(a) Comparison of backpropagation and feedback alignment. Backpropagation is the standard training method for artificial neural networks but is not considered biologically plausible due to the weight transport problem, which requires a symmetric backward pathway for transmitting error signals. Feedback alignment, on the other hand, is a biologically plausible alternative that does not rely on weight transport, instead using a random matrix B to replace the backward synaptic pathway. (b) Test loss during random noise pretraining and subsequent data training. Blue indicates networks pretrained with random noise, and orange indicates networks trained solely with data (without random noise pretraining). (c) Test accuracy during subsequent data training. (d) Histogram of predictive confidence, with vertical lines indicating the averaged accuracy and confidence. (e) The difference between the averaged confidence and accuracy of predictions (w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$; w/o vs. zero, Wilcoxon signed-rank test, NS, $P<10^{-3}$; w/ vs. zero, Wilcoxon signed-rank test, NS, $P=0.752$). (f) Reliability diagram showing expected sample accuracy as a function of binned confidence. (g) Expected Calibration Error (w/o vs. w/ random pretraining, $n_{\text{net}}=10$, Wilcoxon rank-sum test, $P<10^{-3}$).
Supplementary Fig. 4: Pre-calibration of uncertainty in toy-model network.
@@
Supplementary Fig. 5: Pre-calibration of uncertainty in high-dimensional network.
@@
Supplementary Fig. 6: Calibration of unseen real-world data through random noise pretraining.
(a-b) Confidence reduction for various unseen real data during random noise pretraining. (a) CIFAR-10 and SVHN datasets presented to the network during random noise training, with measured accuracy and confidence (Left). Confidence and accuracy for the CIFAR-10 dataset (Right). (b) Confidence distribution for CIFAR-10 and SVHN datasets before and after random noise pretraining. (c) Expected Calibration Error (ECE) for CIFAR-10 and SVHN datasets before and after random noise pretraining.